Below are the probabilities for a set of outcomes. In order to make sure the entire sample space is represented, what must the probability for Outcome 5 be?

Outcome 1 = 0.10

Outcome 2 = 0.30

Outcome 3 = 0.15

Outcome 4 = 0.25

Outcome 5 =
Answer #1

Answer:

0.20

EXPLANATION:

Total Probability = 1.

So, we write:

Outcome 1 + Outcome 2 + Outcome 3 + Outcome 4 + Outcome 5 = 1

Substituting values, we get:

0.10 + 0.30 + 0.15 + 0.25 +Outcome 5 = 1

i.e.,

0.80 + Outcome 5 = 1

So,

Outcome 5 = 1 - 0.80 = 0.20